The perimeter of a parallelogram is 150 cm. One of its side is greater than the other by
25 cm. Find the length of all sides of the parallelogram.

Answer by cleomenius(959)

2x + 2(x + 25) = 150
2x + 2x + 50 = 150
4x = 100
x = 25
To check,
25 + 25 + 50 + 50 = 150.
Cleomenius.